Spent a zillion years doing this and no luck

Many businesses borrow money during periods of increased business activity to finance inventory and accounts receivable. Hudsons Bay Company (HBC) is Canadas largest department store. Each Christmas, HBC builds up its inventory to meet the needs of Christmas shoppers. A large portion of Christmas sales are on credit. As a result, HBC often collects cash from the sales several months after Christmas. Assume that on November 1, 2017, HBC borrowed $6 million cash from Downtown Bank and signed a promissory note that matures in six months. The interest rate was 7.5 percent payable at maturity. The accounting period ends December 31 Required: 1. Indicate the effects (accounts and amounts) of the (a) issuance of the note on November 1, (b) impact of the adjusting entry on December 31, 2017, and (c) the payment of the note and interest on April 30, 2018, on the accounting equation. (Enter your answers in whole dollars. Enter any decreases to accounts with a minus sign.) Date Assets Liabilities (a) Nov. 1.2017 (b) Dec 31, 2017 (c) Apr 30, 2018
Answer #1

1) Effect

Assets = Liabilities + Stockholder's equity
Nov 1,2017 Cash 6000000 Notes payable 6000000
Dec 31,2017 Interest payable 75000 Interest expense -75000
Apr 30,2018 Cash -6225000 Notes payable -6000000 Interest expense -150000
Interest payable -75000

Dec 31,2017 Interest = 6000000*7.5%*2/12 = $75000

Apr 30,2018 Interested = 6000000*7.5%*4/12 = 150000